OpenStudy (killa):

A 50-gallon barrel of milk is 6% butterfat. How much skim milk should be mixed to make milk that is 3% butterfat?

OpenStudy (shadowfiend):

So first you can find how much butterfat is in the 50 gallon barrel: \[\frac{x}{50} = \frac{6}{100}\] Once you have x, the amount of butterfat, you know there is 50 - x gallons of milk. Then you are trying to calculate a value m such that x / (50 - x + m) = (3 / 100). Does that make sense?

OpenStudy (killa):

yeah, i think i get it now. thanks!

OpenStudy (shadowfiend):

Awesome, good to hear!